OpenStudy (anonymous):

****************MEDALS AND FANS******************************************Lana cuts 2 pieces for every 7 inches from a dowel for a craft project, for a total of 7 pieces. What is the length of the dowel initially?

OpenStudy (anonymous):

2 = 7 4 = 14 6 = 21 7 = ? 8 = 35

OpenStudy (shinalcantara):

since in 7 inches 2 pieces are being cut, this only implies that every piece is 7in/2 = 3.5 inches in length since there are 7 pieces then all you have to do is multilpy 3.5in/pc(7pcs)=?

OpenStudy (anonymous):

Not any answer

OpenStudy (shinalcantara):

*multiply \[7(3 \frac{ 1 }{ 2 }) = ?\]

OpenStudy (anonymous):

21

OpenStudy (shinalcantara):

nope

OpenStudy (anonymous):

umm...\[24 (1/2)\]

OpenStudy (shinalcantara):

changing it into improper fraction you'll have \[7(\frac{ 7 }{ 2 })\]

OpenStudy (shinalcantara):

\[\frac{ 7(7) }{ 2 }=?\]

OpenStudy (anonymous):

\[10(1/2)\]

OpenStudy (shinalcantara):

\[\frac{ 49 }{ 2 } = 24 \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Ohhhhhh okay I get it now thanks